create new tag
view all tags

20130127 Laser Repair - LN

During Endo-san's visit, the laser overheated and he spent most of his time trying to fix the high voltage power supply for slave 1 and the cooling system. It was found that the heat exchanger was particularly old and most likely not performing well. As of 2013, they no longer manufacture these parts any more, although it should be possible to find something compatible from another company. Terunuma-san was able to have it repaired by the ATF2 workers. It then performed significantly better - flow rate 12L/min, 'P1' 450 vs 3L/min & 'P1' 660 before. For note, 'P1' on the external chiller has a limit of 680 and upon hitting this limit, the external chiller stopped working. The interlock for the internal cooling circuit was removed by the previous users of the laser system (so said Endo-san) as it was faulty and kept causing false interlock trips. Therefore, the laser has no temperature interlock and will overheat.

It is therefore CRUCIAL to keep an eye on the temperature every hour or so when using the laser.

Everything appeared to be fine after this repair and the laser worked well once Endo-san left, however, about 3 days afterwards, the same overheating problem appeared again - most likely caused by a freshly dislodged piece of material in the heat exchanger. It was about another 2 days after this that the output from the regen dropped sharply.

The laser build up time increased dramatically and the output was around 3mJ seeded and very unstable -> ie the regen was not working at all. This was investigated. The timing system was checked and during the process changed to be suitable for 3Hz. A single signal generator was used to be externally triggered and create a burst of two pulses. Their duty cycle could only be set to 50% which meant the trigger pulses were extremely long, but only the rising edge matters. The laser was working at 3Hz before because of the 3Hz timing triggers from the ATF2, however, the signal generator train used to generate the 6Hz from 3Hz did not respond to a new trigger until it had completed the previous event, causing the flash lamps to flash at approximately 6Hz but only be in time at 1.56Hz, the previous frequency of the laser. After the update, the laser works just fine at 3Hz.

Endo-san (the Coherent engineer) had specifically said not to make large adjustments to the cavity. Only make very small tweaks to the end mirror and the etalon with the aim of decreasing the build up time. However, no improvement was made with this. Therefore, the obvious thing to do is remove the etalon and start from the most basic cavity possible. In removing the etalon, the build up time reduced by almost 200ns. With only a slight tweak of the cavity, the build up time was optimised as well as the output spatial profile (imaged before spatial filter). The output burst energy (unseeded, no etalon) in this case was approximately 23mJ! The etalon was inspected and found to not only be damaged, but be damaged in two distinct locations.

Etalon Damage

Endo-san had specifically stated that the etalon gets damaged when the cavity is unstable due to overheating, which would seem to indicate that he knew this as it happened when he was here. He was able to move the etalon transversely as the approximately 1cm diameter beam can be moved within the 1" etalon. When the laser overheated a second time, further damage was incurred. The increased attenuation in cavity where the pulse experiences exponential gain clearly reduces the output energy.

Therefore, I tried our old trick of raising the etalon and rotating it to further displace the damage from where the laser would now pass through it. With less attenuation, I expected the performance to be better. The etalon was placed in the cavity and it's angle tuned manually (unseeded) to reduce the build up time as much as possible. This required it to be placed at approximately 10 degrees - quite standard. The build up time was significantly lower now. The output energy was approximately 12mJ - not the best, but perfectly acceptable.

In Situ:

From this point on, only small tweaks to the cavity were required to optimise it. It was not possible to increase the output beyond 12mJ. The output was approximately the same seeded and unseeded - only marginally more (and more stable) seeded. The build up time decreased by the usual 40ns when seeded. As well as the energy the following beam profiles were recorded with the TaperCam just after the isolator immediately after the regen.

Seeded with Etalon Before Tuning:

Unseeded without Etalon:

Seeded with Etalon After Tuning:

The regen output was the realigned through the remainder of the laser system with approximately 80% transmission to the linear amps. At 140us flash lamp delay, the output green was also optimised. The downstream alignment was also recovered. The following images were recorded at the VIP where there was suitable attenuation so as not to damage the camera while imaging the beam at full power.

Final Output - Low Energy:

Final Output - High Energy:

Topic attachments
I Attachment History Action Size Date Who Comment
JPEGjpg 2013_01_27_final_output_full_energy.jpg r1 manage 55.4 K 04 Feb 2013 - 21:58 LaurieNevay Final Output - High Energy
JPEGjpg 2013_01_27_final_output_low_energy.jpg r1 manage 13.0 K 04 Feb 2013 - 21:58 LaurieNevay Final Output - Low Energy
JPEGjpg 2013_01_27_seeded_with_etalon.jpg r1 manage 14.6 K 04 Feb 2013 - 21:57 LaurieNevay Seeded with Etalon Before Tuning
JPEGjpg 2013_01_27_seeded_with_etalon_after_tuning.jpg r1 manage 14.1 K 04 Feb 2013 - 21:57 LaurieNevay Seeded with Etalon After Tuning
JPEGjpg 2013_01_27_unseeded_without_etalon.jpg r1 manage 22.8 K 04 Feb 2013 - 21:56 LaurieNevay Unseeded without Etalon
JPEGjpg IMG_4257.jpg r1 manage 111.9 K 04 Feb 2013 - 21:56 LaurieNevay Etalon Damage
JPEGjpg IMG_4260.jpg r1 manage 217.6 K 04 Feb 2013 - 21:55 LaurieNevay In Situ
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r2 - 04 Feb 2013 - LaurieNevay

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback